Nonannual compounding using a calculator​) Jesse Pinkman is thinking about trading cars. He estimates he will still have to borrow ​$34 comma 000 to pay for his new car. How large will​ Jesse's monthly car loan payment be if he can get a 6​-year ​(72 equal monthly​ payments) car loan from the​ university's credit union at an APR of 7.5 percent compounded​ monthly? ​Jesse's monthly car loan payment will be ​$
Answer #1

Calculate the monthly car loan payment as follows:

X L =PMT(7.5%/12,6*12,-34000) Function Arguments PMT Rate 7.5%/12 Nper 6*12 Pv -34000 = 0.00625 = 72 = -34000 * E = number Ty

Monthly payment is $587.86